# Compression, Impact and Hot Rebound Flows from Coronal Rain Downflows

### v1

@@ 5 Discussion and Conclusion
The significant energy of the impact explains its visibility in all AIA/EUV wavelengths. We also found that about $20\%$ of the clump’s kinetic energy is transferred to the thermal and kinetic energy of the rebound flow, along with a negligible wave energy. Hence, the majority of the energy is radiated away upon impact in the transition region. It is interesting to make the analogy to the ‘precipitation limit’ feedback mechanism observed at the much larger scales of the intracluster medium (ICM). In that field, cool cores containing relatively dense and very cold gas are observed to accrete towards the supermassive black holes at the centres of galaxy clusters (Voit et al., 2015). These dense inflows impart momentum to the accretion disks, which power energetic AGN jets that subsequently heat the ICM and can slow down the cooling and cool core generation. Hence, a delicate balance is achieved between the number of cool core accretions and the temperature of the ICM from the AGN jets. These ‘precipitation limit’ cycles with the cool cores are analogous to the TNE cycles with coronal rain found in the solar corona. In this work, this analogy is further extended since we show observationally for the first time that the rebound flow from coronal rain can have a major impact on the hosting coronal loop, not only through increasing the temperature but also modifying its morphology.
@@
Lastly, we observe a continuous intensity enhancement at the footpoint of the loop hosting the rain in SJI 1400 Å, HRIEUV and AIA 193 Å  during the entire observation sequence, which could correspond to the previously reported steady character of the moss (Antiochos et al., 2003; Brooks et al., 2009). These intensity enhancements are lower amplitude but have similar widths as those observed during impact, (both the small and the larger heating widths of $\approx 1$~Mm and $\approx 4$~Mm, respectively). This suggests a steady footpoint heating component of similar heating scale length, which could be the kind predicted by TNE. We can estimate or place a lower bound to the volumetric heating of this component with the following reasoning. The kinetic energy of coronal rain is very likely close to the total integrated volumetric heating from coronal heating over the duration prior to the rain. Although a proper estimation of this hypothesis awaits with numerical modelling, this can be justified by the fact that the footpoint heating in the low corona is partly lost in the transition region being radiated away, but a large part of it goes into producing a mild chromospheric evaporation that increases the coronal density (leading to scaling laws such as RTVS, Serio et al., 1981). Most of this density collapses into the condensation, whose potential energy is then converted into kinetic energy as it falls. If so, we can write:
$E_{CH}\gtrapprox K_{CR},$ (11)
@@
where $E_{CH}$ denotes the total integrated coronal heating from the beginning of the TNE cycle in which the coronal rain forms, and $K_{CR}$ is the kinetic energy of the rain shower, estimated to be $3.09\times 10^{26}$ erg over its 15 min duration. On the other hand, the coronal heating energy is given by:
@@
where $H_{CH}$ is the unknown, roughly steady volumetric heating rate; $s_{H}$ is the heating scale length, which we assume to be equal to the prevalent small width of $\approx 1$ Mm of the intensity brightening observed in AIA 193 Å, HRIEUV and SJI 1400 Å  at the loop footpoint; the factor of 2 is to account for both footpoints assuming similar heating conditions [^1] Numerical simulations indicate that coronal rain often falls towards the footpoint with lower heating rate due to the lower pressure at the footpoint (Froment et al., 2018); and $\tau_{TNE}$ is the period of the cycle. Unfortunately, this observational sequence is not long enough to observe another coronal rain event in the same loop with HRIEUV. However, looking further back in time with AIA 304 Å  reveals another coronal rain event in the same loop at 22:00UT on October 31st, that is, roughly 2 hrs prior to the studied event. We can take the area $A$ to be set by the width of the upward flow at the footpoint (assuming a circular cross-section). Then, we have:
$H_{CH}\gtrapprox\frac{K_{CR}}{2As_{H}\tau_{TNE}}.$ (13)
Replacing by the numerical values, we get a volumetric heating rate $H_{CH}\gtrapprox 2.57\times 10^{-2}$~erg cm-3 s-1. Recently, Ishigami et al. (2024) analysed spectroscopically several coronal loops with Hinode/EIS and found heating scale heights of $s_{H}=10\pm 4$~Mm with a volumetric heating rate of $10^{-2\pm 0.5}$~erg cm-3 s-1. Here we take a much smaller heating scale height, given by the persistent intensity brightening observed at the foopoint. On the other hand, the volumetric heating rate that we find is that we find very close to that reported. Although several other factors can influence TNE and coronal rain formation, this similarity also suggests that the coronal loops studied in Ishigami et al. (2024) may be in a state of TNE. More generally, our study illustrates how coronal rain observations can be used to infer key properties of coronal heating.
@@ Acknowledgements
Data are courtesy of Solar Orbiter, SDO and IRIS. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. The EUI instrument was built by CSL, IAS, MPS, MSSL/UCL, PMOD/WRC, ROB, LCF/IO with funding from the Belgian Federal Science Policy Office (BELPSO); Centre National d’Études Spatiales (CNES); the UK Space Agency (UKSA); the Deutsche Zentrum für Luftund Raumfahrt e.V. (DLR); and the Swiss Space Office (SSO). The building of EUI was the work of more than 150 individuals during more than 10 years. We gratefully acknowledge all the efforts that have led to a successfully operating instrument. SDO is a mission for NASA’s Living With a Star (LWS) program. IRIS is a NASA small explorer mission developed and operated by LMSAL with mission operations executed at NASA Ames Research Center and major contributions to downlink communications funded by ESA and the Norwegian Space Centre. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #545 (‘Observe Local Think Global: What Solar Observations can Teach us about Multiphase Plasmas across Physical Scales’). All images in this manuscript have been made with Matplotlib on Python (Hunter, 2007). This research used version 7.0 of the SunPy open source software package (The SunPy Community et al., 2020).

### v2

@@ 5 Discussion and Conclusion
The significant energy of the impact explains its visibility in all AIA/EUV wavelengths. We also found that roughly $15\%$ of the clump’s kinetic energy is transferred to the thermal and kinetic energy of the rebound flow, along with a negligible wave energy. Hence, the majority of the energy is radiated away upon impact in the transition region. It is interesting to make the analogy to the ‘precipitation limit’ feedback mechanism observed at the much larger scales of the intracluster medium (ICM). In that field, cool cores containing relatively dense and very cold gas are observed to accrete towards the supermassive black holes at the centres of galaxy clusters (Voit et al., 2015). These dense inflows impart momentum to the accretion disks, which power energetic AGN jets that subsequently heat the ICM and can slow down the cooling and cool core generation. Hence, a delicate balance is achieved between the number of cool core accretions and the temperature of the ICM from the AGN jets. These ‘precipitation limit’ cycles with the cool cores are analogous to the TNE cycles with coronal rain found in the solar corona. In this work, this analogy is further extended since we show observationally that the rebound flow from coronal rain can have a major impact on the hosting coronal loop, not only through increasing the temperature but also modifying its morphology.
@@
Lastly, we observe a continuous intensity enhancement at the footpoint of the loop hosting the rain in SJI 1400 Å, HRIEUV and AIA 193 Å  during the entire observation sequence, which could correspond to the previously reported steady character of the moss (Antiochos et al., 2003; Brooks et al., 2009). These intensity enhancements are lower amplitude but have similar widths as those observed during impact, (especially the broad components with widths between 2 Mm and 10 Mm). This suggests a steady footpoint heating component of similar heating scale length, which could be the kind predicted by TNE-TI. We can estimate or place a lower bound to the volumetric heating of this component with the following reasoning. The kinetic energy of coronal rain is very likely close to the total integrated volumetric heating from coronal heating over the duration prior to the rain. Although a proper estimation of this hypothesis awaits with numerical modelling, this can be justified by the fact that the footpoint heating in the low corona is partly lost in the transition region being radiated away, but a large part of it goes into producing a mild chromospheric evaporation that increases the coronal density (leading to scaling laws such as RTVS, Serio et al., 1981). Most of this density collapses into the condensation, whose potential energy is then converted into kinetic energy as it falls. If so, we can write:
$E_{CH}\gtrapprox K_{SH},$ (19)
@@
where $E_{CH}$ denotes the total integrated coronal heating from the beginning of the TNE-TI cycle in which the coronal rain forms, and $K_{SH}$ is the kinetic energy of the rain shower, estimated to be $4.64\times 10^{26}$ erg over its 15 min duration. On the other hand, the coronal heating energy is given by:
@@
where $H_{CH}$ is the unknown, roughly steady volumetric heating rate; $s_{H}$ is the heating scale length, which we assume to be set by the width of the observed persistent broad hot components at the loop footpoint, between 2 Mm and 10 Mm; the factor of 2 is to account for both footpoints assuming similar heating conditions [^1] Numerical simulations indicate that coronal rain often falls towards the footpoint with lower heating rate due to the lower pressure at the footpoint (Froment et al., 2018); and $\tau_{TNE}$ is the period of the cycle. Unfortunately, this observational sequence is not long enough to observe another coronal rain event in the same loop with HRIEUV. However, looking further back in time with AIA 304 Å  reveals another coronal rain event in the same loop at 22:00UT on October 31st, that is, roughly 2 hrs prior to the studied event. We can take the area $A$ to be set by the width of the upward flow at the footpoint (assuming a circular cross-section), and since this area is very similar to that of the rain clumps, we have:
$H_{CH}\gtrapprox\frac{K_{SH}}{2s_{h}A_{SH}\tau_{TNE}}=\frac{\rho_{CR}v_{CR}^{3}}{4s_{H}}\left(\frac{w_{CR}}{w_{SH}}\right)^{2}\frac{\Delta t_{SH}}{\tau_{TNE}},$ (21)
Replacing by the numerical values, we get a volumetric heating rate $H_{CH}\gtrapprox(0.41-2.05)\times 10^{-2}$~erg cm-3 s-1, or $H_{CH}\gtrapprox 10^{-2.04\pm 0.35}$ erg cm-3 s-1. Recently, Ishigami et al. (2024) analysed spectroscopically several coronal loops with Hinode/EIS and found heating scale heights of $s_{H}=10\pm 4$~Mm with a volumetric heating rate of $10^{-2\pm 0.5}$~erg cm-3 s-1. Both, the heating scale height and the volumetric heating rate that we find are very close to those reported. Although several other factors can influence TNE and coronal rain formation, this similarity also suggests that the coronal loops studied in Ishigami et al. (2024) may be in a state of TNE. More generally, our study illustrates how coronal rain observations can be used to infer key properties of coronal heating.
@@ Acknowledgements
We would like to thank Ish Kaul, Peng Oh and Max Gronke for valuable discussions on accretion braking. Similarly, we would like to thank the anonymous referee for the many valuable comments that greatly improved the quality of this manuscript. This research received support by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #545 (‘Observe Local Think Global: What Solar Observations can Teach us about Multiphase Plasmas across Physical Scales’). Data are courtesy of Solar Orbiter, SDO and IRIS. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. The EUI instrument was built by CSL, IAS, MPS, MSSL/UCL, PMOD/WRC, ROB, LCF/IO with funding from the Belgian Federal Science Policy Office (BELPSO); Centre National d’Études Spatiales (CNES); the UK Space Agency (UKSA); the Deutsche Zentrum für Luftund Raumfahrt e.V. (DLR); and the Swiss Space Office (SSO). The building of EUI was the work of more than 150 individuals during more than 10 years. We gratefully acknowledge all the efforts that have led to a successfully operating instrument. SDO is a mission for NASA’s Living With a Star (LWS) program. IRIS is a NASA small explorer mission developed and operated by LMSAL with mission operations executed at NASA Ames Research Center and major contributions to downlink communications funded by ESA and the Norwegian Space Centre. All images in this manuscript have been made with Matplotlib on Python (Hunter, 2007). This research used version 7.0 of the SunPy open source software package (The SunPy Community et al., 2020).